# The complexity of classifying continuous t-norms up to isomorphism

### v2

@@ 1 Introduction
Triangular norms, usually referred to as t-norms [18, 2], are binary operations * on the unit interval $[0,1]$ such that $([0,1],*,1)$ is a commutative ordered monoid. The origins of t-norms can be traced back to the study of probabilistic metric spaces [20, 26], where they were used to generalize the triangle inequality of classical metric spaces. Later, t-norms received wide attention in many-valued logic (see, e.g., [15, 21, 6, 8]), where they provide a mathematical framework for generalizing the concept of conjunction in logic. Beyond logic, t-norms also serve as an indispensable tool in various fields, including game theory [3], decision making [10, 14], non-additive measures [25] and statistics [23].
A t-norm * is called continuous if it is continuous as a function with respect to the usual topology on $[0,1]\times[0,1]$. As the most important class of t-norms, continuous t-norms are a primary focus of research in this field. In particular, the basic fuzzy logic (or BL for short), developed by Hájek [15], is known as a logic of continuous t-norms [15, 5]. It is well known that every continuous t-norm is an ordinal sum of the three basic t-norms [9, 22, 18, 19, 2], i.e, the minimum, the product, and the Łukasiewicz t-norms; we record an equivalent form of this result as Lemma 2.5.
@@ 2 Isomorphisms of continuous t-norms
Given an interval $[a,b]\subseteq\mathbb{R}$, a continuous function $*\colon[a,b]\times[a,b]\to[a,b]$ is called a continuous t-norm on $[a,b]$ [18, 19, 2] if
  • $([a,b],*,b)$ is a commutative monoid, and
@@
For each continuous t-norm $([a,b],*)$ and $q\in[a,b]$, we say that
  • $q$ is idempotent, if $q*q=q$;
@@
(See [18, Proposition 2.3].) Let $([a,b],*)$ be a continuous t-norm. If $q\in[a,b]$ is idempotent, then $p*q=\min\{p,q\}$ for all $p\in(a,b)$.
@@
(See [19, Theorem 2.6].) A continuous t-norm $([a,b],*)$ is isomorphic to the Łukasiewicz t-norm $([0,1],*_{\text{\L}})$ if, and only if, every $q\in(a,b)$ is nilpotent.
@@
Let $([a,b],*)$ be a continuous t-norm. If $([a,b],*)$ is isomorphic to the product t-norm $([0,1],*_{P})$ or the Łukasiewicz t-norm $([0,1],*_{\text{\L}})$, then $p*q<\min\{p,q\}$ for all $p,q\in(a,b)$.
@@
[18, 19, 2] For each continuous t-norm $([0,1],*)$, the set of non-idempotent elements of * in $[0,1]$ is a union of countably many pairwise disjoint open intervals
$\{(a_{\alpha},b_{\alpha})\mid 0<a_{\alpha}<b_{\alpha}<1,\ \alpha\in A,\ A\ \text{is countable}\},$
and for each $\alpha\in A$, the continuous t-norm $([a_{\alpha},b_{\alpha}],*)$ obtained by restricting * to $[a_{\alpha},b_{\alpha}]$ is either isomorphic to the product t-norm $([0,1],*_{P})$ or isomorphic to the Łukasiewicz t-norm $([0,1],*_{\text{\L}})$. Moreover,
$p*q=\min\{p,q\}=p\quad\text{and}\quad q*r=\min\{q,r\}=q$
whenever $0\leqslant p\leqslant a_{\alpha}\leqslant q\leqslant b_{\alpha}\leqslant r\leqslant 1$ for some $\alpha\in A$.
Therefore, it makes sense to denote a continuous t-norm $([0,1],*)$ by
@@
where the intervals $[a_{\alpha},b_{\alpha}]$ $(\alpha\in A)$ are obtained by Lemma 2.5. In order to characterize isomorphisms of continuous t-norms in terms of their interval structures, we define:
@@
  • Ł=$\mathcal{I}_{\text{\L}}^{*}=\{(a_{\alpha},b_{\alpha})\mid\alpha\in A\ \text{and}\ ([a_{\alpha},b_{\alpha}],*)\cong_{t}([0,1],*_{\text{\L}})\}$.
@@
    • for each $\epsilon>0$, there exist non-idempotent elements of * in $(a-\epsilon,a)$ and $(b,b+\epsilon)$.
Let $\mathfrak{S}$ denote the set of all collections of non-empty, pairwise disjoint open subintervals of $[0,1]$ whose union is dense in $[0,1]$. In other words, $\mathcal{S}\in\mathfrak{S}$ if $\mathcal{S}$ is a collection of non-empty, pairwise disjoint open subintervals of $[0,1]$ and
@@
    $\varUpsilon\colon\mathcal{T}\to\mathfrak{S},\quad([a_{\alpha},b_{\alpha}],*)_{\alpha\in A}\mapsto\mathcal{I}_{P}^{*}\cup\mathcal{I}_{\text{\L}}^{*}\cup\mathcal{I}_{M}^{*}.$ (2.iii)
@@
    $\Phi\colon\varUpsilon([0,1],*)\to\varUpsilon([0,1],\bullet)$ (2.iv)
@@
(1$\implies$(2): Let $\varphi\colon([0,1],*)\to([0,1],\bullet)$ be an isomorphism of continuous t-norms. Define
@@
(2$\implies$(1): For each $(a,b)\in\varUpsilon([0,1],*)$, by (2.v) we find an isomorphism of continuous t-norms
@@
Let $2=\{0,1\}$. Write $2^{<\omega}$ for the binary tree, i.e., the set of finite strings made up of 00 and $1$. Recall that a family $J=\{J_{u}\}_{u\in 2^{<\omega}}$ of sets is called a Cantor System (see, e.g., [7, Definition 899]) if for each $u\in 2^{<\omega}$:
@@
  • $J_{u^{\smallfrown}0}$, $J_{u^{\smallfrown}1}\subseteq J_{u}$, where $u^{\smallfrown}0$ and $u^{\smallfrown}1$ refer to the extensions of the finite string $u$ obtained by adding 00 and $1$ at the end of $u$, respectively;
@@ 3 Reducibility of equivalence relations
Note that by [17, Theorem 4.19], the set $\mathcal{C}$ of continuous functions from $[0,1]\times[0,1]$ to $[0,1]$ becomes a Polish space by taking the sup metric
@@
Proof.
By definition, there exist Borel functions $f\colon X\to Y$, $g\colon Y\to X$ such that
$x_{1}Ex_{2}\iff f(x_{1})Ff(x_{2})\quad\text{and}\quad y_{1}Fy_{2}\iff g(y_{1})Fg(y_{2})$
for all $x_{1},x_{2}\in X$, $y_{1},y_{2}\in Y$. Thus, it is easy to see that
$f\times f\colon X\times X\to Y\times Y\quad\text{and}\quad g\times g\colon Y\times Y\to X\times X$
are both Borel functions between Polish spaces.
Suppose that $E=(f\times f)^{-1}(F)$ is a complete analytic subset of $X\times X$. Then $F=(g\times g)^{-1}(E)$ is analytic by [17, Proposition 14.4]. In order to show that $F$ is a complete analytic subset of $Y\times Y$, let $C$ be an analytic subset of a zero-dimensional Polish space $Z$. Since $E$ is complete analytic, there exists a continuous function $h\colon Z\to X\times X$ such that $C=h^{-1}(E)$. Thus $(f\times f)\circ h\colon Z\to Y\times Y$ is a Borel function satisfying $C=((f\times f)\circ h)^{-1}(F)$; since $Y$ is a Polish space and $F$ is analytic, the existence of such a Borel function is enough to guarantee that $F$ is complete analytic (cf. the last paragraph of [17, 26.C]).
Conversely, a similar argument shows that $E$ is complete analytic if so is $F$. ∎
@@
where $<$ is a binary relation symbol and $R_{P},R_{\text{\L}},R_{M}$ are unary relation symbols. By (3.iii) and (3.iv), each $S\in\operatorname{Mod}(L_{1})$ is a quadruple
@@
Let $L_{0}=\{<\}$, where $<$ is a binary relation symbol; let
@@
It is easy to check that $\operatorname{LO}$ is a closed subspace of $\operatorname{Mod}(L_{0})$, and thus a Polish space by Lemma 3.9. For $<^{S_{1}},<^{S_{2}}\in\operatorname{LO}$, it is clear that $<^{S_{1}}\cong_{L_{0}}<^{S_{2}}$ if and only if $(\omega,<^{S_{1}})$ and $(\omega,<^{S_{2}})$ are order isomorphic. Considering the restriction of the equivalence relation $\cong_{L_{0}}$ on $\operatorname{LO}$, the following result is well known:
@@ 4 The complexity of the isomorphism relation between continuous t-norms
Enumerate elements of $\mathbb{Q}\cap[0,1]$ in a non-repeating sequence $\{q_{n}\}_{n\in\omega}$. For each continuous t-norm $([0,1],*)$, define
@@
      $\forall l\in\omega\setminus\{0\}$, $(q_{n})_{*}^{(l)}$ is non-idempotent, and
@@
Conversely, if $q_{n}$ is non-idempotent, then there exists $(x,y)\in\mathcal{I}_{P}^{*}\cup\mathcal{I}_{\text{\L}}^{*}$ such that $q_{n}\in(x,y)$ (by Lemma 2.5). Note that $q_{n}$ is a non-nilpotent element of $([x,y],*)$ by (1)(b); otherwise, if $(q_{n})_{*}^{(l)}=x$ for some $l\in\omega\setminus\{0\}$, then $(q_{n})_{*}^{(l)}$ is idempotent, which violates (1)(b). Thus it follows from Lemma 2.3 that $(x,y)\not\in\mathcal{I}_{\text{\L}}^{*}$, and consequently $(x,y)\in\mathcal{I}_{P}^{*}$. Finally, by (1)(c) and Lemma 2.4 we have $q_{i}\not\in(x,y)$ for all $i<n$. Hence $n=n_{xy}^{*}\in R_{P}^{S_{*}}$.
@@
Conversely, from (3)(a) and (3)(b) we see that every $q\in\mathbb{Q}\cap((q_{l},q_{n})\cup(q_{n},q_{l}))$ is idempotent, and thus, by Lemma 2.5, $(q_{l},q_{n})\cup(q_{n},q_{l})$ is a nonempty interval of idempotent elements. Therefore, there exists $(x,y)\in\mathcal{I}_{M}^{*}$ such that $[q_{l},q_{n}]\cup[q_{n},q_{l}]\subseteq[x,y]$. Finally, (3)(c) guarantees $n_{xy}^{*}=n$, because there exists a non-idempotent element $q_{j}\in(q_{i},q_{n})\cup(q_{n},q_{i})$ for all $i<n$, which means that $q_{i}\not\in[x,y]$. Thus $n\in R_{M}^{S_{*}}$.
@@
Step 1. $\Theta$ is a reduction. Given continuous t-norms $([0,1],*)$ and $([0,1],\bullet)$, we verify that
@@
[…]
@@
for all $x,y\in[a_{n}^{\lessdot},b_{n}^{\lessdot}]$. For any $n\in\omega$, it is easy to see that $([a_{n}^{\lessdot},b_{n}^{\lessdot}],*^{\lessdot})$ is isomorphic to the product t-norm $([0,1],*_{P})$, and it is completely determined by the interval $I_{n}^{\lessdot}=(a_{n}^{\lessdot},b_{n}^{\lessdot})$; that is, for any $*^{\lessdot_{1}},*^{\lessdot_{2}}\in\operatorname{LO}$,
@@
Our main result now arises from the combinations of Propositions 4.16 and 4.17:

### v3

@@ 1 Introduction
Triangular norms, usually referred to as t-norms [18, 2], are binary operations * on the unit interval $[0,1]$ such that ([0,1],,1)([0,1],*,1) is a commutative ordered monoid. The origins of t-norms can be traced back to the study of probabilistic metric spaces [20, 26], where they were used to generalize the triangle inequality of classical metric spaces. Later, t-norms received wide attention in many-valued logic (see, e.g., [15, 21, 6, 8]), where they provide a mathematical framework for generalizing the concept of conjunction in logic. Beyond logic, t-norms also serve as an indispensable tool in various fields, including game theory [3], decision making [10, 14], non-additive measures [25] and statistics [23].
A t-norm * is called continuous if it is continuous as a function with respect to the usual topology on $0,1]\times[0,1]$. As the most important class of t-norms, continuous t-norms are a primary focus of research in this field. In particular, the basic fuzzy logic (or BL for short), developed by Hájek [15], is known as a logic of continuous t-norms [15, 5]. It is well known that every continuous t-norm is an ordinal sum of the three basic t-norms [9, 22, 18, 19, 2], i.e, the minimum, the product, and the Łukasiewicz t-norms; we record an equivalent form of this result as Lemma 2.5.
@@ 2 Isomorphisms of continuous t-norms
Given an interval $a,b]\subseteq\mathbb{R}$, a continuous function $\colon[a,b]\times[a,b]\to[a,b]$ is called a continuous t-norm on $[a,b]$ [18, 19, 2] if
  • ([a,b],,b)([a,b],*,b) is a commutative monoid, and
@@
For each continuous t-norm ([a,b],)([a,b],*) and $q\in[a,b]$, we say that
  • $q$ is idempotent, if qq=qq*q=q;
@@
(See [18, Proposition 2.3].) Let ([a,b],)([a,b],*) be a continuous t-norm. If $q\in[a,b]$ is idempotent, then $p*q=\min\{p,q\}$ for all $p\in(a,b)$.
@@
(See [19, Theorem 2.6].) A continuous t-norm ([a,b],)([a,b],*) is isomorphic to the Łukasiewicz t-norm $([0,1],*_{\text{\L }})$ if, and only if, every $q\in(a,b)$ is nilpotent.
@@
Let ([a,b],)([a,b],*) be a continuous t-norm. If ([a,b],)([a,b],*) is isomorphic to the product t-norm $([0,1],*_{P})$ or the Łukasiewicz t-norm $([0,1],*_{\text{\L }})$, then $p*q<\min\{p,q\}$ for all $p,q\in(a,b)$.
@@
[18, 19, 2] For each continuous t-norm ([0,1],)([0,1],*), there exists a countable index set $A$ and a family of pairwise disjoint nonempty open intervals $\{(a_{\alpha},b_{\alpha}):\alpha\in A\}$ such that:
  • the set of non-idempotent elements is exactly $\bigcup\limits_{\alpha\in A}(a_{\alpha},b_{\alpha})$;
  • for each $\alpha\in A$, the continuous t-norm $([a_{\alpha},b_{\alpha}],*)$ obtained by restricting * to $a_{\alpha},b_{\alpha}]$ is either isomorphic to the product t-norm $([0,1],*_{P})$ or isomorphic to the Łukasiewicz t-norm $([0,1],*_{\text{\L }})$;
  • for any $x,y\in[0,1]$, if there is no $\alpha\in A$ such that $\{x,y\}\subset[a_{\alpha},b_{\alpha}]$, then $x*y=\min(x,y)$.
In particular, if ([0,1],)([0,1],*) is the minimum t-norm $([0,1],*_{M})$, then $A=\varnothing$.
Therefore, it makes sense to denote a continuous t-norm ([0,1],)([0,1],*) by
@@
where the intervals $a_{\alpha},b_{\alpha}]$ $(\alpha\in A)$ are obtained by Lemma 2.5. Specifically, the minimum t-norm $([0,1],*_{M})$ is denoted by $([a_{\alpha},b_{\alpha}],*_{M})_{\alpha\in\varnothing}$. In order to characterize isomorphisms of continuous t-norms in terms of their interval structures, we define:
@@
  • $\mathcal{I}_{\text{\L }}^{*}=\{(a_{\alpha},b_{\alpha})\mid\alpha\in A\ \text{and}\ ([a_{\alpha},b_{\alpha}],*)\cong_{t}([0,1],*_{\text{\L }})\}$.
@@
    • for each $\epsilon>0$ with $a-\epsilon\geqslant 0$, there exist non-idempotent elements of * in $(a-\epsilon,a)$, and for each $\epsilon>0$ with $b+\epsilon\leqslant 1$, there exist non-idempotent elements of * in $(b,b+\epsilon)$.
Let $\mathfrak{S}$ denote the set of all collections of nonempty, pairwise disjoint open subintervals of $[0,1]$ whose union is dense in $[0,1]$. In other words, $\mathcal{S}\in\mathfrak{S}$ if $\mathcal{S}$ is a collection of nonempty, pairwise disjoint open subintervals of $[0,1]$ and
@@
    $\varUpsilon\colon\mathcal{T}\to\mathfrak{S},\quad([0,1],*)=([a_{\alpha},b_{\alpha}],*)_{\alpha\in A}\mapsto\mathcal{I}_{P}^{*}\cup\mathcal{I}_{\text{\L }}^{*}\cup\mathcal{I}_{M}^{*}.$ (2.iii)
    Note that $\varUpsilon([0,1],*)=\mathcal{I}_{P}^{*}\cup\mathcal{I}_{\text{\L }}^{*}\cup\mathcal{I}_{M}^{*}$ is never empty by Lemma 2.5. In particular, for the minimum t-norm $([0,1],*_{M})$, we have $\mathcal{I}_{P}^{*_{M}}=\mathcal{I}_{\text{\L }}^{*_{M}}=\varnothing$ and $\mathcal{I}_{M}^{*_{M}}=\{(0,1)\}$, and consequently $\varUpsilon([0,1],*_{M})=\{(0,1)\}$.
@@
    $\Phi\colon(\varUpsilon([0,1],*),\prec)\to(\varUpsilon([0,1],\bullet),\prec)$ (2.iv)
@@
[…]
@@
[…]
@@
Let $2=\{0,1\}$. Write $2^{<\omega}$ for the binary tree, i.e., the set of finite strings made up of 0 and $1$. Recall that a family $J=\{J_{u}\}_{u\in 2^{<\omega}}$ of sets is called a Cantor System (see, e.g., [7, Definition 899]) if for each $u\in 2^{<\omega}$:
@@
  • $J_{u^{\smallfrown}0}$, $J_{u^{\smallfrown}1}\subseteq J_{u}$, where $u^{\smallfrown}0$ and $u^{\smallfrown}1$ refer to the extensions of the finite string $u$ obtained by adding 0 and $1$ at the end of $u$, respectively;
@@ 3 Reducibility of equivalence relations
Note that by [17, Theorem 4.19], the set $\mathcal{C}$ of continuous functions from $0,1]\times[0,1]$ to $[0,1]$ becomes a Polish space by taking the sup metric
@@
The following lemma will be used later, which states that the property of being a complete analytic set is invariant under Borel bireducibility. The proof is elementary and is left to the reader.
@@
where << is a binary relation symbol and $R_{P},R_{\text{\L }},R_{M}$ are unary relation symbols. By (3.iii) and (3.iv), each $S\in\operatorname{Mod}(L_{1})$ is a quadruple
@@
Let $L_{0}=\{<\}$, where << is a binary relation symbol; let
@@
It is easy to check that $\operatorname{LO}$ is a closed subspace of $\operatorname{Mod}(L_{0})$, and thus a Polish space by Lemma 3.9. For $^{S_{1}},<^{S_{2}}\in\operatorname{LO}$, it is clear that $^{S_{1}}\cong_{L_{0}}<^{S_{2}}$ if and only if $(\omega,<^{S_{1}})$ and $(\omega,<^{S_{2}})$ are order isomorphic. Considering the restriction of the equivalence relation $\cong_{L_{0}}$ on $\operatorname{LO}$, the following result is well known:
@@ 4 The complexity of the isomorphism relation between continuous t-norms
Enumerate elements of $\mathbb{Q}\cap[0,1]$ in a non-repeating sequence $\{q_{n}\}_{n\in\omega}$. For each continuous t-norm ([0,1],)([0,1],*), define
@@
      $\forall l\in\omega\setminus\{0\}$, $(q_{n})_{*}^{(l)}:=\underbrace{q_{n}*q_{n}*\dots*q_{n}}_{l\ \text{times}}$ is non-idempotent, and
@@
Conversely, if $q_{n}$ is non-idempotent, then there exists $(x,y)\in\mathcal{I}_{P}^{*}\cup\mathcal{I}_{\text{\L }}^{*}$ such that $q_{n}\in(x,y)$ (by Lemma 2.5). Note that $q_{n}$ is a non-nilpotent element of ([x,y],)([x,y],*) by (1)(b); otherwise, if $(q_{n})_{*}^{(l)}=x$ for some $l\in\omega\setminus\{0\}$, then $(q_{n})_{*}^{(l)}$ is idempotent, which violates (1)(b). Thus it follows from Lemma 2.3 that $(x,y)\not\in\mathcal{I}_{\text{\L }}^{*}$, and consequently $(x,y)\in\mathcal{I}_{P}^{*}$. Finally, by (1)(c) and Lemma 2.4 we have $q_{i}\not\in(x,y)$ for all $i<n$. Hence $n=n_{xy}^{*}\in R_{P}^{S_{*}}$.
@@
Conversely, from (3)(a) and (3)(b) we see that every $q\in\mathbb{Q}\cap((q_{l},q_{n})\cup(q_{n},q_{l}))$ is idempotent, and thus, by Lemma 2.5, $(q_{l},q_{n})\cup(q_{n},q_{l})$ is a nonempty interval of idempotent elements. Therefore, there exists $(x,y)\in\mathcal{I}_{M}^{*}$ such that $q_{l},q_{n}]\cup[q_{n},q_{l}]\subseteq[x,y]$. Finally, (3)(c) guarantees $n_{xy}^{*}=n$, because there exists a non-idempotent element $q_{j}\in(q_{i},q_{n})\cup(q_{n},q_{i})$ for all $i<n$, which means that $q_{i}\not\in[x,y]$. Thus $n\in R_{M}^{S_{*}}$.
@@
Step 1. $\Theta$ is a reduction. Given continuous t-norms ([0,1],)([0,1],*) and $([0,1],\bullet)$, we verify that
@@
Since $\cong_{L_{1}}$ is analytic, the following corollary is a direct consequence of Proposition 4.15 and [17, Proposition 14.4]:
Corollary 4.16.
$\cong_{t}$ is an analytic equivalence relation on $\mathcal{T}$.
@@
    $\min\{a_{n}^{\lessdot},\ 1-b_{n}^{\lessdot},\ a_{n}^{\lessdot}-b_{m}^{\lessdot}\}\geqslant\dfrac{1}{3^{n+1}}$ whenever $m\lessdot n$.
@@
for all $x,y\in[a_{n}^{\lessdot},b_{n}^{\lessdot}]$. For any $n\in\omega$, it is easy to see that $([a_{n}^{\lessdot},b_{n}^{\lessdot}],*^{\lessdot})$ is isomorphic to the product t-norm $([0,1],*_{P})$, and it is completely determined by the interval $I_{n}^{\lessdot}=(a_{n}^{\lessdot},b_{n}^{\lessdot})$; that is, for any $^{\lessdot_{1}},*^{\lessdot_{2}}\in\operatorname{LO}$,
@@
Our main result now arises from the combinations of Propositions 4.17 and 4.18: